Suppose you are an expert on the fashion industry and wish to gather information to compare the amount earned per month by models featuring Liz Claiborne attire with those of Calvin Klein. The following is the amount ($000) earned per month by a sample of 15 Claiborne models:

$5.6 $6.3 $6.2 $3.2 $7.0 $6.8 $4.0 $3.7 $3.4 $6.1 4.1 4.8 6.5 3.1 4.0

The following is the amount ($000) earned by a sample of 12 Klein models.

$2.9 $3.9 $4.8 $6.2 $1.8 $2.6 $3.5 $3.6 $3.7 $3.1 2.1 2.0

Question: State the decision rule for 0.01 significance level: H0: μClaiborne ≤ μCalvin Klein ; H1: μ Claiborne > μ
Answer #1

This is a right tailed test.

Degrees of freedom = n1 + n2 - 2 = 15 + 12 - 2 = 25

Level of significance = 0.01

Hence,

Critical t value = 2.485

Decision rule: Reject the null hypothesis (Ho) if t > 2.485